If 9^y = x, which of the following represents 9^y - 9^(y-3) :
A) 1/ x^3
B) x^3
C) 729x
D) 729/728 x
E) 728/729 x

Found 2 solutions by Theo, solver91311:
Answer by Theo(13342)

your expression is 9^y - 9^(y-3)
this is equivalent to 9^y - 9^y/(9^3)
since 9^3 = 729, this becomes:
9^y - 9^y/729
since 9^y = x, this becomes:
x - x/729
put everything under a common denominator and you get:
(729*x - x)/729 which becomes:
728*x / 729 which can be shown as:
728/729 * x.

this looks like selection E.

to confirm, let y be any value.
i chose 5.
when y = 5, x = 9^5 = 59049
728/729 * x = 58968
9^5 - 9^(5-3) = 9^5 - 9^2 = 58968
they're the same, confirming the solution is correct.

go with selection E.







Answer by solver91311(24713)





So



You can do the rest of the arithmetic.


John

My calculator said it, I believe it, that settles it


I > Ø